OpenStudy (anonymous):

Solve the following system of equations. x + 3y + 2z = –16 2x – y + 2z = –15 2x – 2y – z = 1

OpenStudy (anonymous):

(2x – y + 2z) - (x + 3y + 2z) = -15 + 16 x - 4y = 1 x = 1+4y (2x – 2y – z) - (2x – y + 2z) = 1 + 15 -y - 3z = 16 2(1+4y) - 2y - z = 2 + 6y - z = 1 6y - z = -1 18y - 3z = -3 -19y = 19, y = -1 x = -3 z = -5 --- (–3, –1, –5)